OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 3 10 4 16 5 28 f(x) = 3(2)x − 2 + 4 f(x) = 3(2)x − 2 − 4 f(x) = one third(2)x − 2 + 4 f(x) = one third(2)x − 2 − 4

OpenStudy (anonymous):

@halorazer @ash2326

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@ganeshie8 plsss help meee

OpenStudy (anonymous):

@Rina.r

ganeshie8 (ganeshie8):

test each option

ganeshie8 (ganeshie8):

lets test option A :- \(\large f(x) = 3*(2)^{x − 2} + 4\)

ganeshie8 (ganeshie8):

test below values in that function. and see if EVERY pair satisfies it. x f(x) 3 10 4 16 5 28

ganeshie8 (ganeshie8):

put x = 3 in the function, do u get 10 ? \(\large f(x) = 3*(2)^{3 − 2} + 4 = 3*(2)^1 + 4 = 6 + 4 = 10 ~~\color{green}{\checkmark}\) put x = 4 in the function, do u get 16 ? \(\large f(x) = 3*(2)^{4 − 2} + 4 = 3*(2)^2 + 4 = 12 + 4 = 16 ~~\color{green}{\checkmark}\) put x = 5 in the function, do u get 28 ? \(\large f(x) = 3*(2)^{5 − 2} + 4 = 3*(2)^3 + 4 = 24 + 4 = 28~~\color{green}{\checkmark}\)

ganeshie8 (ganeshie8):

so option A is the correct funciton that models the given table of data.

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

sorry i had went back to the test i jst came back thankyou i understand!